OpenStudy (anonymous):

Cu(s) + NO3- (aq) ---> Cu2+ (aq) + NO oxidation numbers

OpenStudy (abmon98):

The oxidation state of an uncombined element is zero. Cu is 0 NO3- The more electronegative element in a substance is given a negative oxidation state. Oxygen is the more electronegative and is usally a -2 charge atom. There are 3 oxygen atoms present -2*3=-6 The oxidation state of all ions present in a compound must equal to the charge it carries -6+5=-1 +5 is the oxidation state of N present in NO3- Cu+2 is given and for NO oxygen is usually -2 and the compound is neutral (oxidation state =0) so nitrogen must equal to +2